Calculate the  ΔG° using Hess Law for the reaction: Cu2S(S) + O2(g) ( Cu(s) SO2(g)

2 Cu(s) + S(s)  ( Cu2S(L) Δ    Go  = -86.2 kJ

S(s)  +  O2(g)  (  SO2(g) Δ        Go = -300.1 kJ



1
Expert's answer
2022-04-17T11:04:32-0400

Cu2S(s) + O2(g) → 2Cu(s) + SO2(g), ΔG°x = ???

eq = equation

(1st eq): 2Cu(s) + S(s) → Cu2S(s), ΔG°1 = −86.2 kJ

(2nd eq): S(s) + O2(g) → SO2(g), ΔG°2 = −300.1 kJ


Solution:

According to Hess's law, the heat of reaction depends upon Initial and final conditions of reactants and does not depend of the intermediate path of the reaction.


1) Modify the two given equations to get the target equation:

(1st eq): flip it so as to put Cu2S(s) on the reactant side.

(2nd eq): do nothing. We need one mole of SO2(g) on the product side and that's what we have.


2) Rewrite the two equations with the changes made:

(1*st eq): Cu2S(s) → 2Cu(s) + S(s), ΔG°1* = +86.2 kJ

(2nd eq): S(s) + O2(g) → SO2(g), ΔG°2 = −300.1 kJ


Thus, adding modified equations and canceling out the common species on both sides, we get:

Cu2S(s) + O2(g) → 2Cu(s) + SO2(g)


3) Add the ΔG° values of (1*stand (2ndequations to get your answer:

ΔG°x = ΔG°1* + ΔG°2 = +86.2 kJ + (−300.1 kJ) = −213.9 kJ

ΔG°x = −213.9 kJ


Answer: The ΔG° for the reaction is −213.9 kJ